OpenStudy (anonymous):

Find the cube of 12. A. 36 B. 144 C. 1728 D. 20,736

OpenStudy (anonymous):

@lolacole12

OpenStudy (anonymous):

@Study312

OpenStudy (calculusxy):

\(12^3 = 12 \times 12 \times 12\)

OpenStudy (anonymous):

=1728

OpenStudy (calculusxy):

cube means to multiply the base three times similarly, square means to multiply the base two times

OpenStudy (study312):

cubed is just the number repeated 3 times so \[12^{3}\] is 12 x 12 x 12 = 1,728. what calculusxy said

OpenStudy (anonymous):

ok thanks